OpenStudy (chrisplusian):

can you write x(25-x^2)^1/2 as x(5-x), or (5x-x^2??????

jimthompson5910 (jim_thompson5910):

No, but you can write \[\Large x\left(25-x^2\right)^{\frac{1}{2}}\] as \[\Large x\sqrt{25-x^2}\]

OpenStudy (chrisplusian):

so that isn't the same as\[x((25)^1/2 -(x^2)^1/2 \] ?????

jimthompson5910 (jim_thompson5910):

no it's not

jimthompson5910 (jim_thompson5910):

you can't distribute that exponent through like that

OpenStudy (chrisplusian):

your right I don't know what I am thinking. I have been out f college for eight months and I am forgetting the basic rules for some reason

jimthompson5910 (jim_thompson5910):

that's alright, happens to me all the time